OpenStudy (anonymous):

Calculate the energy for one mole of photons (an einstein) for light absorbed at 6.90 × 102 nm.

OpenStudy (anonymous):

2.455 X 10^5 Joules/Mole Maybe you ccould help answer my question ?

OpenStudy (***[isuru]***):

Hi, u have to solve this step by step... first u have to calculate the energy of a photon and then u have to multiply it by Avogadro constant and find the energy for a mole to find the energy of a photon u can use \[E = h \mu \] where E = energy h = planks constant mu = frequency of the wave

OpenStudy (***[isuru]***):

Now we know that we have to find out frequency of the wave.... but the question has only given the wavelength ... which is 6.90 x 10^2 nm hence all e.m waves have the same velocity in the vacuum , which is the speed of light we can use \[c = \mu \lambda \] where c = veloctiy - speed of light - 3 x 10^8 m/s mu = frequency lambda = wavelength now make mu the subject of the equation \[\mu = \frac{ c }{ \lambda }\] now u can substitute the values we know \[\mu = \frac{ 3 \times 10^{8} }{ 6.9 \times 10 ^{2}\times 10^{-9} }\] the wave length = 6.9 x 10^2 nm = 6.9x 10^2 x 10^(-9) solve it and u will get the value for frequency as \[\frac{ 1 }{2.3 }\times 10 ^{15}\]

OpenStudy (anonymous):

hey, im taking a look at what you have replied but that isnt' the answer. It has to be in J. it's an online assignment. I put these as this answer ( 1.73 x 10 ^ -4 ) and then it got this as response Incorrect. Check to make sure all of the unit conversions have been done properly. Wavelength is given in nanometers, but the speed of light is in meters. [ (6.626 X 10^-34) * (2.998 X 10^8) / (6.90 x 10 ^2) ) ] * 6.022 X 10 ^ 23

OpenStudy (***[isuru]***):

now just plug that value for the frequency in the first equation \[E =h \mu \] planks constant = \[h = 6.624\times10^{-34}\] \[E = 6.624 \times 10^{-34} \times \frac{ 1 }{ 2.3 }\times10^{15}\] and ya will get the energy of a photon as \[E = 2.88 \times 10^{-19} J\] and to find the energy for a mole , multiply that value by Avogadro constant ; \[6.022\times10^{22}\] so the final amount energy is \[E = 2.88 \times 10^{-19} \times 6.022\times10^{22}\] and it will be simplify to this \[E = 1.734 \times 10^{5} J\] here's the final answer....

OpenStudy (***[isuru]***):

did ya get it ?

OpenStudy (anonymous):

Yes ! Finally that is the correct answer now i have to take a look at your recent reply. and make sure im able to do it myself. Thank You.

OpenStudy (***[isuru]***):

u r welcome!